OpenStudy (throat):

integral [(25x²-4)^1/2]/x dx

OpenStudy (kira_yamato):

Do you mean integrate the following with respect to x? \[\frac{\sqrt{25x^2 - 4}}{x}\] If this is true, then you can use integration by parts... http://en.wikipedia.org/wiki/Integration_by_parts

OpenStudy (nikvist):

\[\int\frac{\sqrt{25x^2-4}}{x}dx=\int\frac{5\sqrt{x^2-4/25}}{x}dx=5\int\sqrt{1-\left(\frac{2}{5x}\right)^2}dx\]\[\frac{2}{5x}=\cos{y}\quad,\quad x=\frac{2}{5\cos{y}}\quad,\quad dx=\frac{2\sin{y}}{5\cos^2{y}}dy\]\[5\int\sqrt{1-\left(\frac{2}{5x}\right)^2}dx=5\int\sqrt{1-\cos^2{y}}\frac{2\sin{y}}{5\cos^2{y}}dy=\]\[=2\int\frac{\sin^2{y}}{\cos^2{y}}dy=2\int\left(\frac{1}{\cos^2{y}}-1\right)dy=2(\tan{y}-y)+C=\]\[=2\left(\sqrt{\frac{1}{\cos^2{y}}-1}-y\right)+C=\]\[=2\left(\sqrt{\frac{25x^2}{4}-1}-\arccos\frac{2}{5x}\right)+C=\]\[=\sqrt{25x^2-4}-2\arccos\frac{2}{5x}+C\]